# Maximum entropy principle for quantum processes

### v1

@@ Abstract
The maximum entropy principle states that the maximum entropy among all quantum states with a fixed mean energy is achieved only by the thermal state of given mean energy. In this notes, we prove the maximum entropy principle for quantum processes— the entropy of a quantum channel with fixed mean energy is maximum if and only if the channel is absolutely thermalizing channel with the fixed output thermal state of that mean energy. This allows for an alternate approach to describe emergence of the absolute thermalization processes under energy constraints in the observable universe.
I Introduction
Entropy is a primitive information-theoretic quantity that measures randomness or uncertainty in the system. Entropy of a quantum process quantifies the minimum uncertainty in the output of the process conditioned on the reference system Devetak et al. (2006); Gour and Wilde (2021). It has implications on the erasure cost of the logical output of a bipartite unitary circuit when the eraser has access to the ancillary output Badhani et al. (2025). In quantum information theory, quantum channels describe quantum processes that are completely positive, trace-preserving linear maps. The basic randomness and information measures like entropy, conditional entropy, mutual information, and conditional mutual information of quantum channels provide insight to their underlying causal structures and characteristics Das et al. (2024); Badhani et al. (2025). In this short notes, we state and prove the maximum entropy principle for quantum channels which can be deemed as channelized version of the well-known maximum entropy principle for quantum states Jaynes (1980):
@@ I Introduction
The maximum entropy principle for quantum processes states that among all channels with a fixed mean output energy $E$, i.e., the maximum output mean energy of a channel is $E$, the channel entropy is maximum if and only if the channel is absolutely thermalizing channel with a fixed output thermal state with mean energy $E$.
See Fact 1 and Theorem 1 for the formal statements of the principle for the quantum states and channels, respectively.
I.1 Preliminaries
We first set the standard notations, recall standard definitions, and known facts from the literature.
@@ I.1 Preliminaries
We consider separable Hilbert spaces. Let $\operatorname{St}(A)$ and $\operatorname{Pos}(A)$ denote the sets of all density operators and positive semidefinite operators, respectively, defined on the Hilbert space $\mathcal{H}_{A}$. The Hilbert space $\mathcal{H}_{AB}$ of a composite system $AB$ is given as $\mathcal{H}_{AB}:=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. We may simply denote $\mathcal{H}_{A}$ as $A$. The dimension of $A$ is $|A|:=\dim(A)$. $\mathbbm{1}_{A}$ is the identity operator on $A$. $\omega_{A}$ denotes operator on $A$ and for brevity we may denote it as $\omega$. For $\rho\in\operatorname{St}(AB)$, $\rho_{B}=\operatorname{tr}_{A}(\rho_{AB})$ is the reduced state on $B$. A quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is a completely positive, trace-preserving linear map with input on $A^{\prime}$ and output on $A$. For a linear map $\mathcal{N}_{A^{\prime}\to A}$ and $\psi\in\operatorname{St}(RA^{\prime})$, $\mathcal{N}(\psi_{RA^{\prime}})=\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}\to A}(\psi_{RA^{\prime}})$. Let $\operatorname{Ch}(A^{\prime},A)$ denote the set of all quantum channels $\mathcal{N}_{A^{\prime}\to A}$. Let $\mathcal{R}_{A^{\prime}\to A}$ denote the uniformly mixing map, $\mathcal{R}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\operatorname{tr}(\rho_{A^{\prime}})\mathbbm{1}_{A}$ for all $\rho\in\operatorname{Pos}(A^{\prime})$. A replacer channel $\mathcal{R}^{\omega}_{A\to A^{\prime}}$ always outputs a fixed state $\omega_{A}$ for all possible input states $\rho_{A^{\prime}}$, i.e., , $\mathcal{R}^{\omega}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\omega_{A}$ for all $\rho\in\operatorname{St}(A^{\prime})$.
@@
The Choi operator of a linear map $\mathcal{N}_{A^{\prime}\to A}$ is $\Gamma^{\mathcal{N}}_{RA}=\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}\to A}(\Gamma_{RA^{\prime}})$, where $R\simeq A^{\prime}$ and $\operatorname{id}_{R}$ is the identity supermap with input-output $R$ and $\Gamma_{RA^{\prime}}:=\sum_{i,j=0}^{d-1}\ket{ii}\bra{jj}_{RA^{\prime}}$ with $d=\min\{|R|,|A^{\prime}|\}$. For $d=\min\{|R|,|A|\}<+\infty$, $\Phi_{RA}:=\frac{1}{d}\Gamma_{RA}$ is a maximally entangled state and $\Phi^{\mathcal{N}}_{RA}:=\frac{1}{|A|}\Gamma^{\mathcal{N}}_{RA}$ is a state (Choi state of $\mathcal{N}_{A^{\prime}\to A}$) if $\mathcal{N}_{A^{\prime}\to A}$ is a quantum channel. The Choi operator of a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$ is a product operator $\mathbbm{1}_{R}\otimes\omega_{A}$.
@@
The von Neumann entropy, simply referred to as entropy, of a quantum state $\rho_{A}$ is $S(A)_{\rho}:=-\operatorname{tr}[\rho\log\rho]$. The quantum relative entropy between $\rho\in\operatorname{St}(A),\sigma\in\operatorname{Pos}(A)$ is given by $D(\rho\|\sigma):=\operatorname{tr}\left[\rho(\log\rho-\log\sigma)\right]$ if $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$ else it is +$\infty$+ ∞. $D(\rho\|\sigma)=0$ if and only if $\rho=\sigma$. We note that $S(A)_{\rho}=-D(\rho_{A}\|\mathbbm{1}_{A})$. For a quantum state $\rho\in\operatorname{St}(AB)$, the von Neumann conditional entropy $S(A|B)_{\rho}$ quantifies uncertainty of $A$ conditioned on $B$, $S(A|B)_{\rho}=-\inf_{\sigma\in\operatorname{St}(B)}D(\rho_{AB}\|\mathbbm{1}_{A}\otimes\sigma_{B})$, and $S(A|B)_{\rho}=S(AB)_{\rho}-S(B)_{\rho}$ Cerf and Adami (1997).
@@
Let $\widehat{H}_{A}$ denote the Hamiltonian of quantum system $A$. Throughout this notes, we consider the Hamiltonian $\widehat{H}_{A}$ to be bounded. The mean energy of $\rho\in\operatorname{St}(A)$ is given by $\langle\widehat{H}\rangle_{\rho}:=\operatorname{tr}[\widehat{H}\rho]$. For a Hamiltonian $\widehat{H}_{A}$, the thermal state $\gamma^{\beta}$ is given as $\gamma^{\beta}_{A}={\exp(-\beta\widehat{H}_{A})}/{Z^{\beta}_{A}}$ Lenard (1978), where $\beta=(k_{B}T)^{-1}$ is inverse temperature and $Z^{\beta}_{A}:=\operatorname{tr}[\exp(-\beta\widehat{H}_{A})]$. The temperature $T$ (or, $\beta$) is uniquely determined by the mean energy $\langle\widehat{H}\rangle_{\gamma^{\beta}}$ of the thermal state $\gamma^{\beta}$.
Fact 1 (Jaynes (1980)).
@@
Consider a quantum system $A$ with Hamiltonian $\widehat{H}_{A}$. The maximum entropy principle states that
$\max_{\rho\in\operatorname{St}(A):~{}\operatorname{tr}[\widehat{H}\rho]=E}S(\rho)=S(\gamma^{\beta})$ (1)
for
@@
$\gamma^{\beta}:=\frac{\exp(-\beta\widehat{H})}{\operatorname{tr}[\exp(-\beta\widehat{H})]},\quad\langle\widehat{H}\rangle_{\gamma^{\beta}}=E.$ (2)
@@
Entropy of a quantum channel.— The quantum relative entropy between two quantum channels $\mathcal{N}_{A^{\prime}\to A},\mathcal{M}_{A^{\prime}\to A}$ is defined as $D[\mathcal{N}\|\mathcal{M}]=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{M}(\psi_{RA^{\prime}}))$ Leditzky et al. (2018), and it suffices to consider optimization over pure states $\psi_{RA^{\prime}}$ with $R\simeq A^{\prime}$. The (von Neumann) entropy of an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as $S[\mathcal{N}]:=-D[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]=\inf_{{\psi}\in\operatorname{St}(RA^{\prime})}S(A|R)_{\mathcal{N}(\psi)}$ Gour and Wilde (2021)(see also Devetak et al. (2006); Yuan (2019), in particular (Sohail et al., 2025, Definition 6)), suffices to optimize over pure $\psi_{RA^{\prime}}$ states. For an arbitrary replacer channel $\mathcal{R}^{\omega}$, its minimum output entropy $\inf_{\rho\in\operatorname{St}(A^{\prime})}S(\mathcal{N}(\rho))$ Giovannetti et al. (2004); De Palma et al. (2017) is equal to its entropy,
$S[\mathcal{R}^{\omega}]=S(\omega)=\inf_{\rho\in\operatorname{St}(A^{\prime})}S(\mathcal{N}(\rho)).$ (3)
II Main Result
We now give a formal definition of the mean energy of a quantum channel and then prove the maximum entropy principle for quantum channels.
@@ II Main Result
For a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, we have
$S[\mathcal{R}^{\omega}]=S(\omega),\quad\langle\widehat{H}\rangle_{\mathcal{R}^{\omega}}=\langle\widehat{H}\rangle_{{\omega}}.$ (5)
We now state and prove the maximum entropy principle for quantum processes.
Theorem 1 (Maximum entropy principle for quantum processes).
The maximum entropy among all quantum channels $\mathcal{N}_{A^{\prime}\to A}$ with a fixed mean energy $\langle\widehat{H}\rangle_{\mathcal{N}}=E$, is attained if and only if the channel is absolutely thermalizing channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ with $\langle\widehat{H}\rangle_{\gamma^{\beta}}=E$,
$\max_{\begin{subarray}{c}\ \mathcal{N}\in\operatorname{Ch}(A^{\prime},A):\\ \ \langle\widehat{H}\rangle_{\mathcal{N}}=E\end{subarray}}S[\mathcal{N}]=S[\mathcal{T}^{\beta}]=S(\gamma^{\beta}).$ (6)
@@
For a quantum state $\rho_{AB}$, we know that $S(A)_{\rho}\geq S(A|B)_{\rho}$; the inequality is saturated $S(A)_{\rho}=S(A|B)_{\rho}$ if and only if $\rho_{AB}$ is a product state $\rho_{A}\otimes\rho_{B}$. This is because $S(A)_{\rho}=S(A|B)_{\rho}$ if and only if $D(\rho_{AB}\|\rho_{A}\otimes\rho_{B})=0$. For a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, we have Gour and Wilde (2021); Devetak et al. (2006)
@@
The theorem statement above can be generalized to the cases where chemical potentials Haag et al. (1967); Pathria and Beale (2021) are also considered. Fact 1 will get modified, and Theorem 1 will be revised accordingly. We leave formal statement and rigorous proof for future work. We believe our work provides an axiomatic approach based on the maximum entropy principle to describe absolute thermalization and will be of interest to study quantum to classical transitions Popescu et al. (2006); Leverrier and Cerf (2009); Riera et al. (2012); Brandao and Harrow (2013); Mark et al. (2024).
Acknowledgements.
@@
The authors thank Uttam Singh and Himanshu Badhani for discussions. S.D. acknowledges support from the Science and Engineering Research Board (now ANRF), Department of Science and Technology, Government of India, under Grant No. SRG/2023/000217 and the Ministry of Electronics and Information Technology (MeitY), Government of India, under Grant No. 4(3)/2024-ITEA. S.D. also acknowledges support from the National Science Centre, Poland, grant Opus 25, 2023/49/B/ST2/02468. S.D. thanks Harish-Chandra Research Institute, Prayagraj (Allahabad), India and University of Gdańsk, Gdańsk, Poland for the hospitality during his visits.

### v2

@@ Abstract
The maximum entropy principle, as applied to quantum systems, is a fundamental prescript positing that for a quantum system for which we only have partial knowledge, the maximum entropy state consistent with the partial knowledge is a valuable choice as the system’s state. An intriguing result is that in case the only prior knowledge is of a fixed mean energy, the maximum entropy state turns out to be the thermal state, a ubiquitous state in several arenas, especially in statistical mechanics. We extend the consequences of this principle from static quantum states to dynamic quantum processes. We establish that a quantum channel attains maximal entropy under a fixed mean energy constraint if and only if it is an absolutely thermalizing channel whose fixed output is the thermal state of the same mean energy. This provides an alternative approach for understanding the emergence of absolute thermalization processes within the observable part of the universe under physically natural energy constraints. Our results have potential implications for understanding the informational and thermodynamic utility of quantum channels under physical constraints. As an application, we examine the consequences for private randomness distillation from energy-constrained quantum processes.
Introduction.— The informational aspects of quantum systems and processes provide essential insights into both foundational principles of quantum theory and architect of emerging technologies [1, 2, 3, 4, 5]. Their significance is further underscored by the role of information theory in linking quantum mechanics with thermodynamics, enabling the study of energetics of quantum systems through thermo-informational frameworks [6, 7, 8, 9]. A central concept in these frameworks is entropy, which quantifies the degree of randomness or uncertainty inherent in a system [10]. The maximum entropy principle, introduced by Jaynes [11, 12], has emerged as a widely used tool in the study of thermo-informational aspects of quantum systems [13, 14, 15, 16]. It serves as a foundational guideline for selecting the most unbiased probability distributions consistent with known constraints, making it a natural choice for analyzing quantum systems where often the information is incomplete [17, 18, 19, 20].
A core and long-standing question in quantum physics is why and how thermalization emerges in quantum systems (see e.g., [21, 22, 23, 24, 25, 26, 27]), despite the fundamentally reversible nature of quantum theory due to unitary evolution of closed systems. Given the physical assumption that the mean energy of the observable part of the universe is fixed by nature, and in the absence of any additional prior knowledge about the quantum process under observation, what is the appropriate way to describe the behavior of such a process? Quantum processes are the source of information carriers as well as transformers. They are formally described by completely positive, trace-preserving linear maps, also called quantum channels. To this end, we provide a general extension—or channelized analogue—of the well-known assertion of the maximum entropy principle for quantum states [11, 26] (along with Fact 1). We prove the following key result, thereby exhibiting occurrence of the absolutely thermalization process as a consequence of the maximum entropy principle.
@@
Among all channels constrained by a maximum mean output energy $E$, the channel’s entropy is maximized if and only if it is an absolutely thermalizing channel whose fixed output is a thermal state of the given mean energy $E$.
Quantum processes within the observable part of the universe appear to naturally tend towards absolute thermalization, a phenomenon consistent with the maximum entropy principle. Given that the only initial information known about these quantum processes is their fixed mean energy, it suggests that quantum systems inherently thermalize over time. This means that no matter how a quantum system starts, it will evolve to a state where its energy is distributed as evenly as possible, which is best described by processes that maximize entropy under the constraint of a fixed mean energy.
The entropy of a quantum channel captures the minimum uncertainty in its output when conditioned on a reference system [28, 29]. It determines the thermodynamic work cost of erasing the logical output of a bipartite unitary operation, especially when ancillary outputs are accessible to the eraser [30]. Given that physical transformations of quantum states are describable as quantum channel, the thermo-informational aspects of quantum channels determine the energetics and informational utility of quantum processors and devices [31, 32, 30]. Quantum channels encompass quantum gates (circuits), measurements, and states (density operators) [33, 34]. Our work can have implications for the dynamical resource theory [5, 35] of thermodynamics [7, 36, 37] where the thermodynamic resourcefulness of a quantum channel is determined by its distinguishability from the absolutely thermalizing channel [38].
We begin by presenting the standard notations, definitions, and foundational results as the Preliminaries, which are essential for formulating, proving, and analyzing the implications of our main result (Theorem 1). We then establish our main result and discuss key consequences of the maximum entropy principle for quantum channels, and explore their implications on the private randomness capacity [39] of quantum channels. Finally, we conclude by summarizing our main findings and outlining directions for future research.
@@
Preliminaries.— We consider separable Hilbert spaces. Let $\operatorname{St}(A)$ and $\operatorname{Pos}(A)$ denote the sets of all density operators and positive semidefinite operators, respectively, defined on the Hilbert space $\mathcal{H}_{A}$. The Hilbert space $\mathcal{H}_{AB}$ of a composite system $AB$ is given as $\mathcal{H}_{AB}:=\mathcal{H}_{A}\otimes\mathcal{H}_{B}$. We may simply denote $\mathcal{H}_{A}$ as $A$. The dimension of $A$ is $|A|:=\dim(A)$, which can be finite or infinite. $\mathbbm{1}_{A}$ is the identity operator on $A$. $\omega_{A}$ denotes operator on $A$ and for brevity we may denote it as $\omega$. For $\rho\in\operatorname{St}(AB)$, $\rho_{B}=\operatorname{tr}_{A}(\rho_{AB})$ is the reduced state on $B$. A quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is a completely positive, trace-preserving linear map with input on $A^{\prime}$ and output on $A$. For a linear map $\mathcal{N}_{A^{\prime}\to A}$ and $\psi\in\operatorname{St}(RA^{\prime})$, $\mathcal{N}(\psi_{RA^{\prime}})=\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}\to A}(\psi_{RA^{\prime}})$. Let $\operatorname{Ch}(A^{\prime},A)$ denote the set of all quantum channels $\mathcal{N}_{A^{\prime}\to A}$. Let $\mathcal{R}^{\mathbbm{1}}_{A^{\prime}\to A}$ denote the uniformly mixing map, $\mathcal{R}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\operatorname{tr}(\rho_{A^{\prime}})\mathbbm{1}_{A}$ for all $\rho\in\operatorname{Pos}(A^{\prime})$. A replacer channel $\mathcal{R}^{\omega}_{A\to A^{\prime}}$ always outputs a fixed state $\omega_{A}$ for all possible input states $\rho_{A^{\prime}}$, i.e., $\mathcal{R}^{\omega}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\omega_{A}$ for all $\rho\in\operatorname{St}(A^{\prime})$.
@@
The Choi operator of a linear map $\mathcal{N}_{A^{\prime}\to A}$ is $\Gamma^{\mathcal{N}}_{RA}=\operatorname{id}_{R}\otimes\mathcal{N}_{A^{\prime}\to A}(\Gamma_{RA^{\prime}})$, where $R\simeq A^{\prime}$, $\operatorname{id}_{R}$ is the identity channel from $R\to R$, $\Gamma_{RA^{\prime}}:=\sum_{i,j=0}^{d-1}\ket{ii}\bra{jj}_{RA^{\prime}}$ with $d=\min\{|R|,|A^{\prime}|\}$. The Choi operator of a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$ is a product operator $\mathbbm{1}_{R}\otimes\omega_{A}$.
@@
The von Neumann entropy, simply referred to as entropy, of a quantum state $\rho_{A}$ is $S(A)_{\rho}:=S(\rho_{A}):=-\operatorname{tr}[\rho\log\rho]$. The quantum relative entropy between $\rho\in\operatorname{St}(A),\sigma\in\operatorname{Pos}(A)$ is given by $D(\rho\|\sigma):=\operatorname{tr}\left[\rho(\log\rho-\log\sigma)\right]$ if $\operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma)$ else it is +$\infty$+ ∞. For $\rho,\sigma\in\operatorname{Pos}(A)$, we have $D(\rho\|\sigma)\geq 0$ whenever $\operatorname{tr}[\rho]\geq\operatorname{tr}[\sigma]$, and $D(\rho\|\sigma)=0$ if and only if $\rho=\sigma$. $S(A)_{\rho}=-D(\rho_{A}\|\mathbbm{1}_{A})$. For a quantum state $\rho_{AB}$, the von Neumann conditional entropy $S(A|B)_{\rho}$ quantifies uncertainty of $A$ conditioned on $B$, $S(A|B)_{\rho}=-\inf_{\sigma\in\operatorname{St}(B)}D(\rho_{AB}\|\mathbbm{1}_{A}\otimes\sigma_{B})$, and $S(A|B)_{\rho}=S(AB)_{\rho}-S(B)_{\rho}$ [40] (also see [41]).
@@
Let $\widehat{H}_{A}$ denote the Hamiltonian of a quantum system $A$. The mean energy of $\rho\in\operatorname{St}(A)$ is given by $\langle\widehat{H}\rangle_{\rho}:=\operatorname{tr}[\widehat{H}\rho]$. For a Hamiltonian $\widehat{H}_{A}$, the thermal state $\gamma^{\beta}$ (also called Gibbs state) is given as $\gamma^{\beta}_{A}={\exp(-\beta\widehat{H}_{A})}/{Z^{\beta}_{A}}$ [42], where $\beta:=(k_{B}T)^{-1}$ is the inverse temperature and $Z^{\beta}_{A}:=\operatorname{tr}[\exp(-\beta\widehat{H}_{A})]$. The temperature $T$ (or, $\beta$) is uniquely determined by the mean energy $\langle\widehat{H}\rangle_{\gamma^{\beta}}$ of the thermal state $\gamma^{\beta}$. We have $S(\gamma^{\beta})=\beta\langle\widehat{H}\rangle_{\gamma^{\beta}}+\log Z^{\beta}$.
Fact 1 ([11]).
@@
Consider a quantum system $A$ with Hamiltonian $\widehat{H}_{A}$. The maximum entropy among all states with a given mean energy is attainable if and only if the state is the thermal (Gibbs) state of given mean energy,
$\max_{\rho\in\operatorname{St}(A):\leavevmode\nobreak\ \operatorname{tr}[\widehat{H}\rho]=E}S(\rho)=S(\gamma^{\beta}),$ (1)
@@
for $\gamma^{\beta}:=\frac{\exp(-\beta\widehat{H})}{\operatorname{tr}\left[\exp(-\beta\widehat{H})\right]},\leavevmode\nobreak\ \langle\widehat{H}\rangle_{\gamma^{\beta}}=E$.
@@
Entropy of a quantum channel: The quantum relative entropy between two completely positive linear maps $\mathcal{N}_{A^{\prime}\to A},\mathcal{M}_{A^{\prime}\to A}$ is defined as $D[\mathcal{N}\|\mathcal{M}]=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D(\mathcal{N}(\psi_{RA^{\prime}})\|\mathcal{M}(\psi_{RA^{\prime}}))$ [43, 38, 10], and it suffices to consider optimization over pure states $\psi_{RA^{\prime}}$ with $R\simeq A^{\prime}$. The (von Neumann) entropy of an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as $S[\mathcal{N}]:=-D[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]=\inf_{{\psi}\in\operatorname{St}(RA^{\prime})}S(A|R)_{\mathcal{N}(\psi)}$ [29](see also [28, 44], in particular [38, Definition 6]), suffices to optimize over pure $\psi_{RA^{\prime}}$ states. Taking quantum channel $\mathcal{N}_{A^{\prime}\to A}$ to be a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, its minimum output entropy $\inf_{\rho\in\operatorname{St}(A^{\prime})}S(\mathcal{N}(\rho))$ [45, 46] is equal to its entropy, $S[\mathcal{R}^{\omega}]=S(\omega)=\inf_{\rho\in\operatorname{St}(A^{\prime})}S(\mathcal{N}(\rho))$.
Main Result.— The maximum entropy principle is a fundamental concept in quantum statistical mechanics, asserting that when our knowledge of a quantum system is incomplete, the most reasonable description of its state is the one that maximizes entropy while remaining consistent with prior knowledge of the state [11]. A crucial implication of this principle is that if our only prior knowledge about a quantum system is its fixed mean energy, then the maximum entropy state is precisely the thermal state (Fact 1). Here we give a formal definition of the mean energy of a quantum channel (cf. [47]) and then prove and discuss key consequences of the maximum entropy principle for quantum channels.
@@
For a replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$, we have $S[\mathcal{R}^{\omega}]=S(\omega),\leavevmode\nobreak\ \langle\widehat{H}\rangle_{\mathcal{R}^{\omega}}=\langle\widehat{H}\rangle_{{\omega}}$.
Let $\mathcal{T}^{\beta(E)}_{A^{\prime}\to A}$ denote an absolutely thermalizing channel $\mathcal{T}^{\gamma^{\beta}}_{A^{\prime}\to A}$ of the mean energy $E$, i.e., $\langle\widehat{H}\rangle_{\mathcal{T}^{\beta(E)}}=\langle\widehat{H}\rangle_{{\gamma^{\beta(E)}}}=E$. We first state a key implication of the maximum entropy principle for quantum processes. We then prove that the quantum channel that maximizes the entropy of quantum channels with fixed maximum mean output energy is absolutely thermalizing.
Maximum entropy principle for quantum processes: The maximum entropy principle suggests that the most accurate description of an ongoing quantum process is one that maximizes entropy while remaining consistent with any prior knowledge we have about it. If we assume our only prior knowledge about quantum processes in the observable part of the universe is that they have a fixed mean energy, then, according to this principle, these processes are best characterized as absolutely thermalizing channels. This means that no matter the initial configuration, such a channel will inevitably drive any input quantum state toward a fixed output: a thermal state consistent with the process’s mean energy. In essence, based on the following theorem, every initial quantum state will naturally evolve to this thermal equilibrium because the underlying quantum channel inherently leads to thermalization.
Theorem 1.
The maximum entropy among all quantum channels $\mathcal{N}_{A^{\prime}\to A}$ with a fixed mean energy $\langle\widehat{H}\rangle_{\mathcal{N}}=E$, is attained if, and only if, the channel is absolutely thermalizing $\mathcal{T}^{\beta(E)}_{A^{\prime}\to A}$, where $\langle\widehat{H}\rangle_{\mathcal{T}^{\beta(E)}}=E$,
$\max_{\begin{subarray}{c}\ \mathcal{N}\in\operatorname{Ch}(A^{\prime},A):\\ \ \langle\widehat{H}\rangle_{\mathcal{N}}=E\end{subarray}}S[\mathcal{N}]=S[\mathcal{T}^{\beta(E)}].$ (3)
@@
For a quantum state $\rho_{AB}$, we know that $S(A)_{\rho}\geq S(A|B)_{\rho}$; the inequality is saturated $S(A)_{\rho}=S(A|B)_{\rho}$ if and only if $\rho_{AB}$ is a product state $\rho_{A}\otimes\rho_{B}$. This is because $S(A)_{\rho}=S(A|B)_{\rho}$ if and only if $D(\rho_{AB}\|\rho_{A}\otimes\rho_{B})=0$. For a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, we have [29, 28]
@@
The theorem statement above can be generalized to the cases where chemical potentials [48, 26] are also considered. Fact 1 will get modified by considering the generalized thermal (Gibbs) state, and Theorem 1 will be revised accordingly. We leave formal statement and rigorous proof for future work.
[…]
Private or intrinsic randomness: There is a trade-off relation between the entropy $S[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, for $|A|<\infty$, and its private randomness capacity $P_{\rm random}[\mathcal{N}]$ [39][10, Eq. (128)]
$S[\mathcal{N}]+P_{\rm random}[\mathcal{N}]=\log|A|.$ (6)
The private randomness capacity of a channel $\mathcal{N}_{A^{\prime}\to A}$ is the maximum rate, in an asymptotic setting of i.i.d. uses of the channel, at which the receiver accessing output $A$ can extract private randomness, against an eavesdropper accessing extension $E$ of an isometric channel extension $\mathcal{V}^{\mathcal{N}}_{A^{\prime}\to AE}$ of the channel $\mathcal{N}$, when the sender sends states through the channel. Under the physically natural assumption of the energy constraint, the private randomness capacity is minimal for the absolutely thermalizing channel. The minimum private randomness capacity among all quantum channels $\mathcal{N}_{A^{\prime}\to A}$ with a fixed mean energy $\langle\widehat{H}\rangle_{\mathcal{N}}=E$, is attained if, and only if, the channel is absolutely thermalizing $\mathcal{T}^{\beta(E)}_{A^{\prime}\to A}$,
$\min_{\begin{subarray}{c}\ \mathcal{N}\in\operatorname{Ch}(A^{\prime},A):\\ \ \langle\widehat{H}\rangle_{\mathcal{N}}=E\end{subarray}}P_{\rm random}[\mathcal{N}]=P_{\rm random}[\mathcal{T}^{\beta(E)}].$ (7)
The trade-off relation (6) resembles closely with the trade-off relation between the entropy $S(\rho_{A})$ of a quantum state $\rho_{A}$ and its objective information $I(\rho_{A}):=D(\rho_{A}\|\pi_{A})$ [52, 53, 54, 51, 55], where $\pi_{A}:=\mathbbm{1}/|A|$ is the maximally mixed state and $|A|<\infty$,
$S(\rho)+I(\rho)=\log|A|.$ (8)
The objective information of a state is related to the amount of extractable work [56, 57] and intrinsic randomness [58] in the state. We can express the private randomness capacity $P_{\rm random}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ as $P_{\rm random}[\mathcal{N}]=D[\mathcal{N}\|\mathcal{R}^{\pi}]$, where $\mathcal{R}^{\pi}_{A^{\prime}\to A}$ is the replacer channel which always outputs the maximally mixed state $\pi_{A}$. The entropic quantity $D[\mathcal{N}\|\mathcal{R}^{\pi}]$ can be deemed as a measure of the intrinsic randomness of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ [10, Section IV.A.], given that $D(\rho_{A}\|\pi_{A})$ is a measure of the intrinsic randomness of a state $\rho_{A}$. Trade-off relations (8) and (6) are state–channel analogue of each other. For an arbitrary replacer channel $\mathcal{R}^{\omega}_{A^{\prime}\to A}$ we have $S[\mathcal{R}^{\omega}]=S(\omega)$ and $P_{\rm random}[\mathcal{R}^{\omega}]=I(\omega)$. We can arguably also interpret the absolute thermalization as a consequence of the minimum intrinsic randomness principle, a quantum information-theoretic dual to the maximum entropy principle.
Discussion.— Understanding the thermo-informational aspects of quantum processes is crucial for both advancing quantum technologies and deepening our foundational knowledge of quantum mechanics. Information-theoretic quantities like entropy, conditional entropy, mutual information, and conditional mutual information are vital for discerning the causal structure and unique properties of these underlying quantum dynamics [10, 30].
In this work, we have utilized the maximum entropy principle–a robust framework for drawing unbiased inferences from incomplete data–and applied it specifically to quantum channels. This provides an axiomatic approach to understanding absolute thermalization [59, 13, 15, 16]. Our findings, which are based on the maximum entropy principle, offer new insights into the research area of quantum-to-classical transitions [6, 60, 61, 62].
Furthermore, we have developed an argument for evaluating the thermodynamic and informational resourcefulness of a quantum channel under energy constraints. This can be achieved by quantifying how much a given channel deviates from an absolutely thermalizing channel [38]. This approach holds promising implications for thermodynamic resource theories [7, 37] related to quantum channels.
Note added: This topic is also studied in the parallel work by Philippe Faist and Sumeet Khatri, titled “The thermal channel: maximum channel entropy principle and the microcanonical channel" [63].
@@
Acknowledgments.— The authors thank Himanshu Badhani, Karol Horodecki, and Uttam Singh for discussions. S.D. acknowledges support from the SERB (now ANRF), Department of Science and Technology, Government of India, under grant no. SRG/2023/000217 and the Ministry of Electronics and Information Technology (MeitY), Government of India, under grant no. 4(3)/2024-ITEA, and the National Science Centre, Poland, under grant Opus 25, 2023/49/B/ST2/02468. S.D. thanks Harish-Chandra Research Institute, Prayagraj (Allahabad), India and University of Gdańsk, Gdańsk, Poland for the hospitality during his visits.